# The Open Journal of Astrophysics

### v2

@@ Abstract
[…]
@@ 1 Introduction
Weak lensing, a subtle yet rich effect that maps the distribution of dark matter and measures cosmic structure growth, is a key cosmological probe for the Legacy Survey of Space and Time (LSST, Ivezić et al. (2019)) of the Vera C. Rubin Observatory, as well as for the Roman (Akeson et al., 2019) and Euclid (Scaramella et al., 2022) missions. These next-generation surveys will be incredibly powerful, achieving sub-percent levels of statistical precision. The intrinsic alignment (IA) of galaxies, a result of their interactions with large-scale structure and other galaxies, can contaminate weak lensing measurements, necessitating accurate and efficient IA modeling in preparation for future analyses (e.g. Troxel & Ishak, 2015; Krause et al., 2016; Blazek et al., 2019; Fortuna et al., 2021; Hoffmann et al., 2022; Secco et al., 2022; Campos et al., 2023; Samuroff et al., 2024; Paopiamsap et al., 2024).
@@
Intrinsic alignment (IA) modeling has traditionally relied on analytic approaches, such as perturbation theory (e.g. Hirata & Seljak, 2004; Bridle & King, 2007; Blazek et al., 2015, 2019; Vlah et al., 2020, 2021; Maion et al., 2023; Bakx et al., 2023; Chen & Kokron, 2023). However, these analytic models often struggle to accurately capture nonlinear effects. In cosmology, simulation-based approaches that account for both gravitational and baryonic effects, spanning scales from sub-parsec to gigaparsec, have provided profound insights into cosmological evolution (e.g. Villaescusa-Navarro et al., 2021; Nelson et al., 2021; Pillepich et al., 2018; Delgado et al., 2023). These methods have the potential to better capture IA effects compared to purely analytic models. Magnetohydrodynamic simulations, often referred to as “hydro” simulations, incorporate baryonic effects but exhibit significant variance in their predictions depending on the simulation suite and the sub-grid physics models employed at sub-parsec scales. This variance includes disagreement on IA effects measured in different simulation suites (e.g. Tenneti et al., 2016; Samuroff et al., 2021). Additionally, these simulations are computationally expensive. Given these challenges, gravity-only N-body simulations, such as AbacusSummit (Maksimova et al., 2021) and Quijote (Villaescusa-Navarro et al., 2020), provide a cost-effective and general alternative. These simulations avoid the need to specify sub-parsec-scale baryonic physics, but inherently lack galaxy formation and evolution processes. To bridge this gap, various halo occupation distribution (HOD) models have been employed to populate halos from N-body simulations with galaxies. This approach has been extended to include galaxy populations that exhibit correlated alignments (e.g. Joachimi et al., 2013; Hoffmann et al., 2022; Van Alfen et al., 2024).
@@
Despite its utility, HOD modeling can still be computationally demanding, as it requires the generation of extensive galaxy catalogs from halo catalogs and the application of estimators to extract IA and clustering signals. To reduce computational cost and time associated with IA modeling, surrogate modeling based on existing simulations represents a promising avenue of research. One approach that can offer both precision and efficiency is deep learning (DL), by training neural network (NN) surrogates to accelerate numerical simulations. NN’s have seen many different scientific applications, including in cosmology (see Dvorkin et al. (2022) for a review), with the availability of large datasets and powerful GPU-driven computation. They not only provide new insights, but also have the potential to accelerate numerous analyses when deployed on GPUs. The benefits of NN-based surrogate models are not exclusive to forward modeling, as the differentiability of such models can also be exploited in accelerating inverse problems using differentiable sampling techniques.
@@
In this work, we introduce a NN-based approach to emulating HOD simulations that models both galaxy bias and IA statistics, referred to as IAEmu. This is the first attempt at directly modeling IA statistics from HOD simulations using NNs. IAEmu directly models galaxy position and shape correlations for an HOD simulation, bypassing both the generation of a full galaxy catalog and the simulation step itself. This approach offers a significant speed advantage over traditional HOD-based modeling. Additionally, IAEmu successfully models galaxy shape statistics, whose stochasticity is dominated by galaxy shape noise, as discussed in Van Alfen et al. (2024). IAEmu successfully captures the mean behavior of these noisier statistics, which would otherwise require multiple realizations of the underlying HOD. It also estimates galaxy shape noise (aleatoric uncertainty) and quantifies its own epistemic uncertainty – reflecting uncertainty in the predicted correlation amplitudes – primarily due to limited training data. IAEmu’s uncertainty estimates enable one to assess the reliability of these predictions and further enable error propagation in modeling pipelines that incorporate IAEmu. We further show the benefits of accelerated parameter inference (i.e., inverse problems) using gradient-based sampling techniques with IAEmu, exploiting the fact that NNs are differentiable models.
@@
Related Work. Several previous works have constructed simulation-based emulators for cosmological statistics, with a focus on matter or galaxy density. Zhai et al. (2019) constructed Gaussian process-based emulators based on the AEMULUS Project’s N-body simulations for nonlinear galaxy clustering. Kwan et al. (2023) similarly used a Gaussian process-based emulator, HOD modeling, and the Mira-Titan Suite of N-body simulations to predict galaxy correlation functions, building on earlier work from the same group (Lawrence et al., 2010). The BACCO simulation project (Aricò et al. 2021a, Aricò et al. 2021b) built NN emulators to include nonlinear and baryonic effects from simulations. These projects emulate various cosmological statistics from simulations, but do not include IA. Jagvaral et al. (2022), Jagvaral et al. (2024), and Jagvaral et al. (2023b) developed generative models trained on the TNG100 simulation (Nelson et al., 2021) to emulate IA in hydrodynamic simulations, but these models do not emulate statistics. Our work is the first to emulate galaxy-IA correlation statistics using simulated galaxy catalogs.
@@
Paper Organization. This paper is organized as follows. Section 2 provides a background on the HOD simulation and correlation function estimators, as well as the procedure for generating and cleaning the training and test data. Section 3 introduces the IAEmu architecture and the process for training IAEmu. In Section 3, we also analyze the generalization performance of IAEmu on held-out data and characterize the quality of predictions based on the predicted aleatoric and epistemic uncertainty. Finally, in Section 4, we validate the out-of-distribution (OOD) performance of IAEmu on a non-HOD-based signal from the TNG300 suite of simulations (Nelson et al., 2015; Pillepich et al., 2017; Springel et al., 2017; Nelson et al., 2017; Naiman et al., 2018; Marinacci et al., 2018) by obtaining a posterior on alignment parameters, which is compared with the HOD-based approach. We summarize our main results in Section 5.
@@
Figure 1: Ranges of HOD parameters used in generating the training data from halotools-IA. We generate uniform random values for the four occupation parameters, excluding $\log M_{\text{min}}$. These values are based on a linear relationship with $\log M_{\text{min}}$, serving as a central line. The range for random values extends $4\cdot\text{RMSE}$ surrounding this line. To clarify the visualization, $\sigma_{\log(M)}$ is displayed separately from other mass variables. Each panel presents published data from Zheng et al. (2007) as a solid line, while the dotted line illustrates the linear fit to $\log M_{\text{min}}$, with the shaded area indicating the range for uniform random value selection for each parameter. Not shown here are the two alignment parameters, $\mu_{\text{cen}}$ and $\mu_{\text{sat}}$, which both vary uniformly on the range $[-1,1]$ with no relation to these five occupation parameters.
@@ 2.1 HOD Background & Estimators
Given a catalog of dark matter halos, we generate a galaxy catalog using an HOD model. This model consists of several interconnected components: (1) an occupation component, which populates halos with galaxies, (2) a phase space component, which determines the spatial distribution of galaxies within halos, and (3) an alignment component, which models galaxy intrinsic alignments. The halotools package constructs these HOD-based galaxy catalogs following this framework. Specifically, it employs the halo model (Cooray & Sheth, 2002; Asgari et al., 2023) along with alignment models introduced in Van Alfen et al. (2024), providing a flexible approach for generating mock galaxy catalogs while simultaneously tracking intrinsic alignments. We refer to this extension of halotools, which incorporates IA information, as halotools-IA. This structure enables the rapid generation of multiple galaxy catalogs using consistent occupation, phase space, and alignment parameters. Depending on the chosen HOD parameters, a given halo may or may not host a central galaxy – the most massive galaxy residing at the halo’s center. Additionally, halos may contain satellite galaxies, which are distributed throughout the halo.
@@
where $n(r)$ is the number of galaxies separated by distance $r$, and $\bar{n}(r)$ is the expected number of galaxies separated by distance $r$ for a random distribution. This equation is simpler than the Landy-Szalay estimator (Landy & Szalay, 1993) and may be suboptimal in some cases (Singh et al., 2017). However, due to the periodic nature of the simulation box, halotools-IA can use analytical randoms, mitigating much of this suboptimality. This estimator is also much faster and is sufficient for our HOD models. The $\omega$ correlation is defined as
@@
and measures how similarly two galaxies at positions $\bf{x}$ and $\bf{x}+\bf{r}$ are oriented. A positive $\eta$ indicates that the orientations tend to be aligned, while a negative value means they tend to be perpendicular. For both $\omega$ and $\eta$, ${\bf x}$ is the position vector of a given galaxy, ${\bf r}$ is the separation vector between two galaxies, $\hat{r}$ is the unit vector of the separation vector ${\bf r}$, and $\hat{e}$ is the galaxy orientation unit vector. The factor of $1/3$ in these equations accounts for the fact that
$\frac{1}{4\pi}\int_{0}^{2\pi}\int_{0}^{\pi}\cos^{2}\theta\sin\theta\,d\theta\,d\phi=\frac{1}{3}\;,$ (4)
@@
Correlation functions are measured for simulated galaxies across 20 bins, evenly spaced in logarithmic scale, between a minimum separation of $0.1\,h^{-1}{\rm Mpc}$ and a maximum separation of $16\,h^{-1}{\rm Mpc}$. In future work, the maximum range of this correlation could be extended. However, for this dataset, we chose this maximum separation because the number of galaxies $n$ increases with $r$, and the computational cost of measuring correlations scales as $\mathcal{O}(n)\log(n)$. In general, galaxies at $r\leq 1\,h^{-1}{\rm Mpc}$ are considered to be in the “1-halo regime” (galaxies within the same halo) and galaxies outside this range are in the “2-halo regime” (galaxies residing in separate halos).
@@ 2.2 Dataset Generation
To train IAEmu, we generate galaxy catalogs using halotools-IA, incorporating seven HOD and IA parameters derived from an existing dark matter halo catalog that is consistent with a realistic cosmology. We use dark matter catalogs from the Bolshoi-Planck (BolPlanck) simulations, which are available directly through halotools-IA for this purpose (Klypin et al., 2011). We populate halos with galaxies following occupation equations from Zheng et al. (2007). To choose physically plausible values for the five occupation parameters used by these two models, we select the best-fit HOD parameter values for the Sloan Digital Sky Survey sample from Table 1 of Zheng et al. (2007). Further details of how we employ these occupation methods as well as a discussion of the phase space and alignment models are given in Appendix A.
@@
The two alignment parameters, $\mu_{\rm cen}$ and $\mu_{\rm sat}$, govern the shape of the Dimroth-Watson distribution from which galaxy misalignments are sampled, as introduced in Van Alfen et al. (2024). More specifically, an alignment parameter value of 0 corresponds to a uniform distribution in $\cos{(\theta)}$, where $\theta$ is the galaxy misalignment angle, indicating randomly oriented galaxies. Values approaching $1$ indicate perfectly aligned galaxies, while values approaching $-1$ correspond to perpendicular alignments.
@@
To generate training data, we generate evenly spaced values of $\log{M_{min}}$ within the range $[11,15]$. For each of these points, we draw a value for each of the other four occupation parameters uniformly from a region ±$\pm 4\cdot{\rm RMSE}$ around the linear fit to $\log{M_{\text{min}}}$, where RMSE refers to the root mean squared error between the observed values of each occupation parameter and their corresponding values predicted by the linear fit. The two alignment parameters, $\mu_{\rm cen}$ and $\mu_{\rm sat}$, are each sampled uniformly on $[-1,1]$. The sampling procedure was designed to yield physically viable configurations for training IAEmu. It was observed that certain input configurations could lead to an absence of galaxies in specific bins, resulting in NaN values in the correlation functions. This issue frequently arises at small scales, or when the values of $\log{M_{\text{min}}}$ are sufficiently large, making the halos that host galaxies rare. To address this, such cases were removed. Additionally, as a further screening measure, we impose a restriction on input configurations that yield $\xi/\xi_{\text{DM}}$ values exceeding 100, as these are deemed unphysical.
@@
With these input parameter values, we generate galaxy catalogs using halotools-IA and measure the three correlations described in Section 2.1. As HOD modeling is inherently stochastic, we generate 10 realizations of a galaxy catalog for each given set of input values for training. The multiple realizations can enable IAEmu to distinguish the signal from the shape noise of the data, and they later serve to quantify the performance of IAEmu for the noisier correlations. As discussed in more detail in Appendix D of Van Alfen et al. (2024), shape noise dominates over sample variance in the HOD models used for $\omega$ and $\eta$ correlations. Thus, we can capture most of the statistical variance by re-aligning galaxies through these extra realizations. The final dataset has $110,526$ parameter choices, with 10 realizations, for a total of $1,105,260$ entries. These are split into a 70% train, 10% validation, and 20% test set with unique input parameters in each subset. The training data was generated using a combination of 2.4 GHz Intel E5-2680 CPUs and 2.1 GHz Intel Xeon Platinum 8176 CPUs. The simulations were parallelized across 150 cores, split evenly to allow simultaneous calculation of the correlation functions.

### v4

@@ Abstract
The intrinsic alignments (IA) of galaxies, a significant contaminant in weak lensing analyses, arise from correlations in galaxy shapes driven by gravitational tidal interactions and galaxy formation processes. Understanding IA is therefore essential for deriving accurate cosmological inferences from weak lensing surveys. However, most IA modeling relies on a combination of perturbative approaches, which cannot describe nonlinear scales, and expensive simulation-based approaches. In this work, we introduce IAEmu, a neural network-based emulator designed to predict the galaxy position-position ($\xi$), position-orientation ($\omega$), and orientation-orientation ($\eta$) correlation functions, and their associated uncertainties, using halo occupation distribution (HOD)-based mock galaxy catalogs. Compared to the simulated catalogs, IAEmu exhibits an approximately $3\%$ average error for $\xi$ and $5\%$ for $\omega$, while capturing the stochasticity in $\eta$, avoiding overfitting this inherently noisier statistic. Importantly, the emulator also provides aleatoric and epistemic uncertainties, which when analyzed jointly, can help identify regions in parameter space where IAEmu’s predictions may be less reliable. Furthermore, we demonstrate the model’s generalization to a non-HOD based signal by fitting alignment parameters from the tng300 hydrodynamical simulations. Since IAEmu is a fully differentiable neural network, it enables approximately a $10{,}000\times$ speed-up in mapping HOD parameters to correlation functions when deployed on a GPU, compared to conventional CPU resources. This substantial acceleration also facilitates solving inverse problems more efficiently by supporting gradient-based sampling algorithms. As such, IAEmu offers an efficient and promising surrogate model for halo-based galaxy bias and IA modeling with the potential to expedite model validation in Stage IV weak lensing surveys.
@@ 1 Introduction
Weak lensing, a subtle yet rich effect that maps the distribution of dark matter and measures cosmic structure growth, is a key cosmological probe for the Legacy Survey of Space and Time (LSST, lsst) of the Vera C. Rubin Observatory, as well as for the Roman (roman) and Euclid (euclid) missions. These next-generation surveys will be incredibly powerful, achieving sub-percent levels of statistical precision. The intrinsic alignment (IA) of galaxies, a result of their interactions with large-scale structure and other galaxies, can contaminate weak lensing measurements, necessitating accurate and efficient IA modeling in preparation for future analyses (e.g. TROXEL20151; krause16; Blazek_2019; fortuna21a; Hoffmann_2022; secco22; campos23; samuroff24; Paopiamsap_2024).
@@
Intrinsic alignment (IA) modeling has traditionally relied on analytic approaches, such as perturbation theory (e.g. Lamman_2024; Bridle_2007; Blazek_2015; Blazek_2019; vlah_2020; vlah_2021; maion_2023; bakx_2023; chen_2023). However, these analytic models often struggle to accurately capture nonlinear effects. In cosmology, simulation-based approaches that account for both gravitational and baryonic effects have provided profound insights into cosmological evolution (e.g. camels; nelson2021illustristng; pillepich18; delgado_2023). These methods have the potential to better capture IA effects compared to purely analytic models. Magnetohydrodynamic simulations, often referred to as “hydro” simulations, incorporate baryonic effects but exhibit significant variance in their predictions depending on the simulation suite and the sub-grid physics models employed at sub-parsec scales. This variance includes disagreement on IA effects measured in different simulation suites (e.g. tenneti16; samuroff21). However, the volumes required for modern cosmological analyses are prohibitively large for hydrodynamical simulations. Given these challenges, gravity-only $N$-body simulations, such as AbacusSummit (abacussummit) and Quijote (quijote), provide a cost-effective and general alternative. These simulations avoid the need to specify sub-parsec-scale baryonic physics, but inherently lack galaxy formation and evolution processes. To bridge this gap, various halo occupation distribution (HOD) models have been employed to populate halos from $N$-body simulations with galaxies. This approach has been extended to include galaxy populations that exhibit correlated alignments (e.g. Joachimi_2013; Hoffmann_2022; vanalfen_2023). At the scales required for cosmological inference, emulators built on these $N$-body and HOD frameworks are essential for rapidly generating accurate mock galaxy catalogs across large parameter spaces.
@@
Despite its utility, HOD modeling can still be computationally demanding, as it requires the generation of extensive galaxy catalogs from halo catalogs and the application of estimators to extract IA and clustering signals. To reduce computational cost and time associated with IA modeling, surrogate modeling and emulation based on existing simulations represents a promising avenue of research. One approach that can offer both precision and efficiency is deep learning (DL), by training neural network (NN) surrogates to accelerate numerical simulations. NNs have seen many different scientific applications, including in cosmology (see dvorkin2022machine for a review), with the availability of large datasets and powerful GPU-driven computation. They not only provide new insights, but also have the potential to accelerate numerous analyses when deployed on GPUs. The benefits of NN-based surrogate models are not exclusive to forward modeling, as the differentiability of such models can also be exploited in accelerating inverse problems using differentiable sampling techniques.
@@
In this work, we introduce an NN-based approach to emulating HOD simulations that models both galaxy bias and IA statistics, referred to as IAEmu. IAEmu is the first NN-based model to directly predict galaxy position and shape correlations from HOD parameters, eliminating the need to rerun simulations or generate explicit galaxy catalogs. This approach offers a significant speed advantage over traditional HOD-based modeling. Additionally, IAEmu successfully models galaxy shape statistics, whose stochasticity is dominated by galaxy shape noise, as discussed in vanalfen_2023. IAEmu successfully captures the mean behavior of these noisier statistics, which would otherwise require multiple realizations of the underlying HOD. It also estimates galaxy shape noise (aleatoric uncertainty) and quantifies its own epistemic uncertainty – reflecting uncertainty in the predicted correlation amplitudes – primarily due to limited training data. IAEmu’s uncertainty estimates enable one to assess the reliability of these predictions and further enable error propagation in modeling pipelines that incorporate IAEmu. We further show the benefits of accelerated parameter inference (i.e., inverse problems) using gradient-based sampling techniques with IAEmu, exploiting the fact that NNs are differentiable models.
@@
Related Work. Several previous works have constructed simulation-based emulators for cosmological statistics, with a focus on matter or galaxy density. Zhai_2019 constructed Gaussian process-based emulators based on the aemulus Project’s $N$-body simulations for nonlinear galaxy clustering. Kwan_2023 similarly used a Gaussian process-based emulator, HOD modeling, and the Mira-Titan Suite of $N$-body simulations to predict galaxy correlation functions, building on earlier work from the same group (2010ApJ...713.1322L). The Bacco simulation project (2021arXiv210414568A, 2021MNRAS.506.4070A) built NN emulators to include nonlinear and baryonic effects from simulations. These projects emulate various cosmological statistics from simulations, but do not include IA. Jagvaral_2022, jagvaral24_arxiv, and 2023arXiv231211707J developed generative models trained on the tng100 simulation (nelson2021illustristng) to emulate IA in hydrodynamic simulations, but these models do not emulate statistics. Our work is the first to emulate galaxy-IA correlation statistics using simulated galaxy catalogs.
@@
Paper Organization. This paper is organized as follows. Section 2 provides a background on the HOD simulation and correlation function estimators, as well as the procedure for generating and cleaning the training and test data. Section 3 introduces the IAEmu architecture and the process for training IAEmu. In Section 3, we also analyze the generalization performance of IAEmu on held-out data and characterize the quality of predictions based on the predicted aleatoric and epistemic uncertainty. Finally, in Section 4, we validate the out-of-distribution (OOD) performance of IAEmu on a non-HOD-based signal from the tng300 suite of simulations (Nelson_2015; Pillepich_2017; Springel_2017; Nelson_2017; Naiman_2018; Marinacci_2018) by obtaining a posterior on alignment parameters, which is compared with the HOD-based approach. We summarize our main results in Section 5.
@@
Figure 1: Ranges of HOD parameters used in generating the training data from halotools-IA. We generate uniform random values for the four occupation parameters, excluding $\log M_{\text{min}}$. These values are based on a linear relationship with $\log M_{\text{min}}$, serving as a central line. The range for random values extends $4\cdot\text{RMSE}$ surrounding this line. To clarify the visualization, $\sigma_{\log(M)}$ is displayed separately from other mass variables. Each panel presents published data from Zheng_2007 as a solid line, while the dotted line of the same color illustrates the linear fit to $\log M_{\text{min}}$, with the shaded area indicating the range for uniform random value selection for each parameter. Not shown here are the two alignment parameters, $\mu_{\text{cen}}$ and $\mu_{\text{sat}}$, which both vary uniformly on the range [1,1][-1,1] with no relation to these five occupation parameters.
@@ 2.1 HOD Background & Estimators
Given a catalog of dark matter halos, we generate a galaxy catalog using an HOD model. This model consists of several interconnected components: (1) an occupation component, which populates halos with galaxies, (2) a phase space component, which determines the spatial distribution of galaxies within halos, and (3) an alignment component, which models galaxy intrinsic alignments. The halotools Hearin_2017 package constructs these HOD-based galaxy catalogs following this framework. Specifically, it employs the halo model (cooray_2002; asgari_2023) along with alignment models introduced in vanalfen_2023, providing a flexible approach for generating mock galaxy catalogs while simultaneously tracking intrinsic alignments. We refer to this extension of halotools, which incorporates IA information, as halotools-IA (VanAlfen2025). This structure enables the rapid generation of multiple galaxy catalogs using consistent occupation, phase space, and alignment parameters. Depending on the chosen HOD parameters, a given halo may or may not host a central galaxy – the most massive galaxy residing at the halo’s center. Additionally, halos may contain satellite galaxies, which are distributed throughout the halo.
@@
where $n(r)$ is the number of galaxies separated by distance $r$, and $\bar{n}(r)$ is the expected number of galaxies separated by distance $r$ for a random distribution. This equation is simpler than the Landy-Szalay estimator (landy_1993) and may be suboptimal in some cases (sukhdeep_2017), since it omits the random-pair corrections that reduce variance and account for survey geometry that are present in Landy-Szalay, making it more sensitive to noise and boundary effects. However, due to the periodic nature of the simulation box, halotools-IA can use analytical randoms, mitigating much of this suboptimality. This estimator is also computationally faster and is sufficient for our HOD models. The $\omega$ correlation is defined as
@@
and measures how similarly two galaxies at positions $\bf{x}$ and $\bf{x}+\bf{r}$ are oriented. A positive $\eta$ indicates that the orientations tend to be aligned, while a negative value means they tend to be perpendicular. For both $\omega$ and $\eta$, $\bf x$} is the position vector of a given galaxy, $\bf r$} is the separation vector between two galaxies, $\hat{r}$ is the unit vector of the separation vector $\bf r$}, and $\hat{e}$ is the galaxy orientation unit vector that specifies the intrinsic orientation of each galaxy’s major axis. The factor of $1/3$ in these equations accounts for the fact that
$\frac{1}{4\pi}\int_{0}^{2\pi}\int_{0}^{\pi}\cos^{2}\theta\sin\theta\,\mathrm{d}\theta\,\mathrm{d}\phi=\frac{1}{3}\;,$ (4)
@@
In this work, correlation functions are measured for simulated galaxies across 20 $r$ bins, evenly spaced in logarithmic scale, between a minimum separation of $0.1\,h^{-1}{\rm Mpc}$ and a maximum separation of $16\,h^{-1}{\rm Mpc}$. In future work, the maximum range of this correlation could be extended. However, for this dataset, we chose this maximum separation because the number of galaxies $n$ increases with $r$, and the computational cost of measuring correlations scales as $\mathcal{O}(n)\log(n)$. In general, galaxies at $r\leq 1\,h^{-1}{\rm Mpc}$ are considered to be in the “1-halo regime” (galaxies within the same halo) and galaxies outside this range are in the “2-halo regime” (galaxies residing in separate halos).
@@ 2.2 Dataset Generation
Table 1: Bolshoi-Planck simulation parameters.
Simulation Particle Mass $\Omega_{m,0}$ $\sigma_{8}$ $n_{s}$ $h$ $L_{\rm box}$ $z$
($h^{-1}M_{\odot}$) ($h^{-1}$Mpc)
Bolshoi-Planck $\sim 10^{8}$ 0.30711 0.82 0.96 0.70 250 0
To train IAEmu, we generate galaxy catalogs using halotools-IA, incorporating seven HOD and IA parameters derived from an existing dark matter halo catalog that is consistent with a realistic cosmology. We use dark matter catalogs from the Bolshoi-Planck simulations, which are available directly through halotools-IA for this purpose (Klypin_2011). The Bolshoi-Planck simulation is a $250\;h^{-1}\text{Mpc}$ box, with correlations in this work measured up to $16\;h^{-1}\text{Mpc}$. Simulation parameters for Bolshoi-Planck can be found in Table 1. We populate halos with galaxies following occupation equations from Zheng_2007. To choose physically plausible values for the five occupation parameters used by these two models, we select the best-fit HOD parameter values for the Sloan Digital Sky Survey (SDSS) sample from Table 1 of Zheng_2007. Further details of how we employ these occupation methods as well as a discussion of the phase space and alignment models are given in Appendix A.
@@
The two alignment parameters, $\mu_{\rm cen}$ and $\mu_{\rm sat}$, govern the shape of the Dimroth-Watson distribution from which galaxy misalignments are sampled, as introduced in vanalfen_2023. More specifically, an alignment parameter value of 0 corresponds to a uniform distribution in $\cos{(\theta)}$, where $\theta$ is the galaxy misalignment angle, indicating randomly oriented galaxies. Values approaching $1$ indicate perfectly aligned galaxies, while values approaching 1-1 correspond to perpendicular alignments.
To generate training data, we generate evenly spaced values of $\log{M_{min}}$ within the range $[11,15]$, covering the typical halo mass scales that host galaxies. For each of these points, we draw a value for each of the other four occupation parameters uniformly from a region $\pm 4\cdot{\rm RMSE}$ around the linear fit to $\log{M_{\text{min}}}$, where RMSE refers to the root mean squared error between the fiducial SDSS values of each occupation parameter from Zheng_2007 and their corresponding values predicted by the linear fit. The two alignment parameters, $\mu_{\rm cen}$ and $\mu_{\rm sat}$, are each sampled uniformly on [1,1][-1,1]. Earlier iterations of the dataset employed Latin hypercube sampling to fully sample the parameter space; however, this frequently resulted in galaxy samples that were unrealistic. For this reason, we restricted our sampling to values lying near the empirical trends of the SDSS-based HOD fits, which ensures that the resulting correlations remain physically plausible while still spanning a representative range of parameter space.
@@
Despite the constrained sampling strategy, it was observed that certain input configurations could lead to an absence of galaxy pairs in specific separation bins, resulting in NaN values in the correlation functions. This issue frequently arises at small scales, or when the values of $\log{M_{\text{min}}}$ are sufficiently large, making the halos that host galaxies rare. To address this, corresponding galaxy catalogs were removed. Additionally, as a further screening measure, we impose a restriction on input configurations that yield $\xi/\xi_{\text{DM}}$ values exceeding 100, as these are deemed unphysical.
@@
With these input parameter values, we generate galaxy catalogs using halotools-IA and measure the three correlations described in Section 2.1. As HOD modeling is inherently stochastic, we generate 10 realizations of a galaxy catalog for each given set of input values for training. The multiple realizations can enable IAEmu to distinguish the signal from the shape noise of the data, and they later serve to quantify the performance of IAEmu for the noisier correlations. We note that the sample variance does also contribute to the variance in the correlations, but it is always subdominant to the shape noise (see Appendix D of vanalfen_2023 and Equation A3 of Chisari_2013). Thus, we can capture most of the statistical variance by re-aligning galaxies through these extra realizations. The final dataset has $110,526$ parameter choices, with 10 realizations, for a total of $1,105,260$ entries. These are split into a 70% train, 10% validation, and 20% test set with unique input parameters in each subset. The training data was generated using a combination of 2.4 GHz Intel E5-2680 CPUs and 2.1 GHz Intel Xeon Platinum 8176 CPUs. The simulations were parallelized across 150 cores, split evenly to allow simultaneous calculation of the correlation functions.